# GASP: Gaussian Splatting for Physic-Based Simulations

### v1

@@ 4 Experiments
This experimental section demonstrates that GASP is a universal method that can be applied to both small objects and large scenes. Furthermore, it allows for the modeling of interactions between objects and breaking them into pieces. GASP is also capable of working with dynamic scenes.
@@
The Shiny Dataset [20] contains examples of static objects with varying textures. In such a case, all Gaussian components are converted to GaMeS-based representations, after which training of the physical engine is conducted. Specifically, GASP is employed to move the these representations in accordance with a dynamic function. As can be observed in Fig. 1 and Fig. 3, the model is capable of producing high-quality animation renders. In Tab. 1(a), we compare the FID score obtained by GS and GaMeS on static scenes to GASP during simulations. As we can see, our simulation obtained a slightly worse score but still with high-quality renders.
Figure 11: Controlling animations using physics engines allows for application to dynamic objects. This figure compares animations of a dynamic object at three different times under the same physical parameters. This experiment uses GASP integrated with the D-MiSo model [21].
@@
We show that our pipeline also works on dynamic scenes using sample objects from the D-NeRF dataset [17]. We use D-MiSo [21] to model dynamic scenes by flat Gaussians. Then we select random frame and use physical engine. Fig. 1 and Fig. 11 present how the objects change according to physical engine and timeline. In Tab. 1(b), we present the FID score in three different time-steps of dynamic scenes. As we can see, our renders obtain similar scores, which means that our model produces consistent simulations.s.
@@
Our model can model the interaction of static objects. In this case, each element of the scene is trained separately. Also, the properties of each component are specified by different parameters. In Fig. 4 we show a sand-based teddy bear and a glass cup. As we can see, the sand falls smoothly into the cup. Let us take the example of interaction between two objects. In Fig. 9 we show how our model handles the formation of artifacts. In Fig. 5 we integrate small duck objects into a full scene. As we can see, the ducks bounce off the table and fall to the ground. Fig. 1 demonstrates the application of physics to two interacting objects: a falling ball and a car reacting to its impact.
Breaking Objects
GASP uses triangle soup to represent flat Gaussians, so we can easily break objects into pieces. In Fig. 12 we show how a cup can be broken. It is important to note that we need a special physics engine to model such a situation. In this example we use additional information about the mesh to train GaMeS on mesh faces (for details see [22]). In this case, Gaussians lie perfectly on the mesh. The initial Gaussian splatting must be trained with additional conditions, but the final GASP modifications are applied analogously to the previous examples (except that here the model is parameterized by mesh).
Figure 12: In this experiment, we create a simulation of a shattering cup using physics simulations. The GaMeS model is used to create a scenario where Gaussians parameters can be adjusted using the mesh.
@@
We demonstrate that physics modeling can function as a black box by using two different simulation approaches.
Taichi__\__elements [^2] https://github.com/taichi-dev/taichi__\__elements is a highly efficient physics engine for the simulation of multi-material continuum mechanics, originally developed in Python using taichi[6], which is designed for high-performance parallel numerical computation. It supports multiple materials, including water, elastic objects, snow, and sand.
Blender [^3] https://www.blender.org version 3.6 is the most popular open source application for a variety of 3D computer graphics tasks. Our GaMeS-based representation can be imported into Blender, where it can be seen as a triangle soup. Using tools such as mesh deformation, we were able to smoothly simulate soft body physics. The simulation objects were then exported and rendered using the GS renderer. Due to the high speed of the latter process, this method allows us to render our simulation faster for realistic scenes than in the case of using traditional 3D methods in Blender. The animations created in Blender utilized the ”Lattice” and ”Mesh Deform” modifiers, with the mesh being defined by the user, such as a sphere or a cube.
Table 1: Frechet Inception Distance (FID) is calculated in two settings: static scenes (Bonsai, Fox) and dynamic scene standup. GASP obtains slightly worse FID concerning static model GS and GaMeS; see Tab. 1(a). GASP gives similar FID across dynamic scenes; see Tab. 1(b).
FID ($\downarrow$)
Method Bonsai Fox
3DGS 19.71 51.04
GaMeS 19.64 47.62
Ours 38.33 88.59
(a) Comparison of FID score obtained by classical GS and GaMeS for static scenes and renders on images taken across GASP simulation.
Frame FID ($\downarrow$)
Standup
Start 133.55
Mid 112.79
End 142.90
(b) FID score for three different times obtained by GASP. As we can see, the simulation is consistent across dynamic scenes.
@@ 5 Conclusion
In this paper, we introduce GASP, a model that integrates high-quality rendering with accurate physical simulation. By leveraging a GaMeS-based representation of objects, our approach enables the simulation of basic phenomena, such as crushing, and more complex scenarios, including cracking and rainfall. This demonstrates the model’s versatility and effectiveness in handling a wide range of physical interactions within simulated environments. Experimental section, shows that GASP work in many difference scenario and produce high quality simulations.
@@
Our model produces artifacts when the simulation changes shapes essentially.

### v3

@@ 4 Experiments
This experimental section demonstrates that GASP is a universal method that can be applied to small objects and large scenes alike. Furthermore, it allows for the modeling of interactions between objects and breaking them into pieces. GASP is also capable of working with dynamic scenes.
$t_{0}$ $t_{i}$ $t_{T}$
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 10: Jelly cube falling simulation created in Blender using soft-body and cloth modifiers to mimic gel-like behavior. In this setup, each cube was trained independently with GaMeS, allowing every cube in the scene to exhibit distinct properties and more realistic collision interactions.
$t_{0}$ $t_{i}$ $t_{T}$
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 11: Flag movement simulation created in Blender using the Lattice modifier, based on a triangle soup.
@@
The Shiny Dataset [20] contains examples of static objects with varying textures. In such a case, all Gaussian components are converted to GaMeS-based representations, after which the physics engine is used to modify their positions. Specifically, GASP is employed to move these representations in accordance with a dynamic function. As can be observed in Fig. 1, Fig. 2, Fig. 9, Fig. 10, and Fig. 11 the model can produce high-quality animation renders.
Figure 12: Controlling animations using physics engines allows for application to dynamic objects. This figure compares animations of a dynamic object at two different times under the same physical parameters. This experiment uses GASP integrated with the D-MiSo model [21].
@@
We show that our pipeline also works with dynamic scenes using sample objects from the D-NeRF dataset [17]. We use D-MiSo [21] to model dynamic scenes using flat Gaussians. Then, we select a frame and use a physics engine. Fig. 1 and Fig. 12 present how the objects change according to the physics engine and timeline.
@@
We show that GASP can model the interaction of static objects, as shown in Fig. 6. In this case, each element of the scene is trained separately. In addition, the properties of each component are specified by different parameters. Fig. 1 shows the application of physics to two interacting objects: a falling ball and a car reacting to its impact. In Fig. 2, we integrate small duck objects into a full scene. As can be seen, the ducks bounce off the table and fall to the ground. In Fig. 7, we show a sand-based teddy bear and a glass cup. As we can see, the sand falls smoothly into the cup. In the same figure, we also show how our model handles the formation of artifacts.
@@
We demonstrate that physics modeling can function as a black box by using three different simulation approaches.
Genesis [^1] https://github.com/Genesis-Embodied-AI/Genesis [2] is a versatile physics simulation platform designed for Robotics, Embodied AI, and Physical AI applications, featuring a robust physics engine, ultra-fast robotics simulation. It supports diverse physics solvers including, but not limited to, MPM, FEM, PBD, as well as various material models. In our work, we focus on the MPM solver.
Taichi_\_elements [^2] https://github.com/taichi-dev/taichi_\_elements is a highly efficient physics engine for the simulation of multi-material continuum mechanics, originally developed in Python using taichi [7], which is designed for high-performance parallel numerical computation. It supports multiple materials, including water, elastic objects, snow, and sand.
We also carried out an experiment to visually and numerically compare our method with PhysGaussian (see Fig. 13). Specifically, we trained a GaMeS model and integrated it directly into their engine. Their approach performs MPM simulations on Gaussian centers followed by covariance calculations. In contrast, we utilized their engine but performed simulations on designated points from Gaussians without additional covariance computations. In Tab. 1 shows that, qualitatively, the FID remains consistent with PhysGaussian. Since we use the same engine and the same number of Gaussians, the result is expected and maintains fairness in comparison. Tab. 2 shows that incorporating hierarchical structuring significantly accelerates the simulation process. Moreover, visual comparison is included in Fig. 13.
Method Vasedeck Fox Teddybear Ficus Bottle
PhysGaussian 68.66 167.16 227.91 86.63 261.48
GASP (our) 68.22 168.55 244.39 83.98 254.84
Table 1: FID score obtained by PhysGaussian [24] and GASP for static scenes and renders on images taken across simulations. For both methods, we use the same physics engine and trained GS model. In the case of GASP, we treated each triangle vertex as a Gaussian and used their positions across the simulation to calculate the actual parameters of each Gaussian.
Method Vasedeck Fox Teddybear Ficus Bottle
PhysGaussian 53.82 110.37 618.07 89.29 46.42
GASP (our) 74.55 168.81 614.55 186.44 87.03
GASP (our with Core) 17.94 64.69 351.18 19.93 18.80
Table 2: Comparison of time in seconds required to finish the simulation for various models between PhysGaussian and GASP. Both methods use the same GS models and physics engine. For GASP, we also show the speed for simulations where only Core Gaussians were used during simulation.
Figure 13: Comparison of renders obtained by GASP and PhysGaussian on ficus and fox simulations. For both methods, the same trained model was used, and the simulations were conducted using the physics engine used in PhysGaussian. For GASP we used vertices of triangle soup as positions used during simulation, after each step their position was used to calculate the parameters of Gaussians represented by each triangle.
$\alpha$ 1.0 2.0 3.0 4.0 5.0 $\infty$
FID 156.99 162.91 168.21 168.73 168.59 188.21
Table 3: Ablation study of the proposed strategy with varying $\alpha$ values. The FID score generally increases with larger $\alpha$, highlighting the effectiveness of the method. $\infty$ indicates the strategy was not applied.
We conduct an ablation study to assess the impact of the proposed Gaussian Control Rule. The results, summarized in Table 3, report the value of $\alpha$ used in Equation 12 (first row) and the corresponding FID score on the simulation shown in Figure 7 (second row). The findings indicate that the strategy yields consistent improvements. Furthermore, the FID score generally increases with larger $\alpha$ values, with only a slight drop observed between $\alpha=4.0$ and $\alpha=5.0$.
@@ 5 Conclusion
In this paper, we introduce GASP, a pipeline that integrates high-quality rendering with accurate physics simulation. Using a GaMeS-based representation of objects, our approach enables the simulation of basic phenomena, such as crushing, and more complex scenarios, including cracking. This demonstrates the model’s versatility and effectiveness in handling a wide range of physics interactions within simulated environments. Conducted experiments show that GASP works in many different scenarios and produces high-quality simulations. We observe that the additional control rules we proposed for hierarchical simulations significantly accelerate the simulation process. This approach resembles the particle sampler commonly utilized in many physics engines.
@@
Our model may produce artifacts when the simulation causes drastic changes in the shape of the original object. Moreover, the GASP pipeline does not account for variations in lighting interaction with the modified surface. Since our method is built upon the GaMeS representation, it also inherits its limitations. Flat Gaussians may require a large number of components to accurately capture highly curved or anisotropic regions, while the mesh-free triangle soup lacks explicit connectivity. On the other hand, in mesh-based settings, the representation quality is constrained by the resolution of the mesh.
6 Acknowledgements
The project “Effective rendering of 3D objects using Gaussian Splatting in an Augmented Reality environment” (FENG.02.02-IP.05-0114/23) is carried out within the First Team programme of the Foundation for Polish Science co-financed by the European Union under the European Funds for Smart Economy 2021-2027 (FENG). The work of P. Spurek was supported by the National Science Centre (Poland), Grant No. 2021/43/B/ST6/01456.